# FineFake: A Knowledge-Enriched Dataset for Fine-Grained Multi-Domain Fake News Detection

### v2

@@ 3.1. Multi-domain News Collection
Given our overarching objective of constructing a multimodal dataset and benchmark exclusively in English, we judiciously filter out text-only posts and news articles published in languages other than English. Subsequently, we implement data deduplication techniques to eradicate redundant news items, thereby mitigating the risk of data leakage. Moreover, to uphold stringent data quality standards, we systematically eliminate entries containing unqualified images or excessively brief textual content.
3.2. Semi-manual Knowledge Alignment
@@
One notable feature of FineFake is its incorporation of pertinent background knowledge, which complements the original multi-modal content. This integration of contextual information and common knowledge holds immense potential in furnishing invaluable signals to advance the identification of fake news. To accomplish the integration of this crucial common knowledge, we adopt a semi-manual knowledge alignment strategy. Specifically, we leverage the entity link tool (ferragina2010tagme, ; ZHANG2024103693, ) to adeptly recognize named entities within the news text. To ensure the accuracy of extracted named entities, we initially set the threshold at 0.3 to compile the initial list of entities and subsequently adjust the threshold to 0.5 to identify entities with higher confidence. Entities extracted within the threshold range of 0.3 to 0.5 undergo further verification by human annotators to ascertain alignment with the contextual meaning, thereby achieving entity disambiguation. Subsequently, we retrieve all triples within Wikidata with a distance of one from the entity, retaining relationship-type triples with other entities and attribute triples that describe information about the entity. The obtained entities and their corresponding information as triples in knowledge graph will be showcased in conjunction with the original multi-modal content. The entities and their corresponding information, represented as triples within the knowledge graph, are then seamlessly showcased in conjunction with the original multi-modal content. This augmentation enriches the dataset by providing an additional layer of contextual understanding, thereby facilitating more nuanced analyses and interpretations of fake news instances.
Figure 2. Basic information and statistic analysis on FineFake
3.3. Fine-grained Human Annotation
@@
Despite news collected from fact-checking websites comes with authenticity labels assigned by professionals, the other two types of news lack reliable category labels. Therefore, we have engaged professional annotators to label these unlabeled news articles. In the verification stage, 5 annotators are asked to annotate the binary classification label and fine-grained label for each data. As data from official media and online stream media are collected based on the extended information declared by fact-checking websites, we initially utilize the corresponding labeled news from fact-checking websites as reliable references for each data instance. Annotators then will annotate the fine-grained labels and binary classification labels with the help of reference from fact-checking websites. To promote consistency and mitigate subjectivity, we implemented inter-annotator agreement measures. Each news will be independently labeled by five annotators, allowing for the assessment of labeling consistency and the calculation of agreement scores (e.g., Cohen’s kappa coefficient). Any discrepancies or disagreements were then resolved through discussions and consensus among the annotators, ensuring a strict and unified labeling approach.
Through this refined categorization scheme, FineFake facilitates a more nuanced examination of the reasons underlying the falseness of news, empowering researchers to develop sophisticated detection methodologies that account for distinct modalities, inconsistencies, and the intersection between textual and visual elements.
3.4. Downstream Fake News Detection Tasks
Based on the constructed FineFake dataset, we propose three downstream tasks to evaluate the performance of state-of-the-art(SOTA) fake news detection models under various scenarios.
Binary classification task. The primary objective in fake news detection models is to accurately classify news articles as either fake or true. Therefore, we employ the conventional approach of binary classification as the evaluation task. To investigate the efficacy of external knowledge in enhancing fake news detection, we conduct two types of binary classification tasks: one without any external knowledge and another with knowledge augmentation. The loss function is defined as follows:
(1) $\mathcal{L}(y,\hat{y})=-y\log(\hat{y})-(1-y)\log(1-\hat{y}),$
@@ 3.4. Downstream Fake News Detection Tasks
Fine-grained classification task. Currently, both fake news detection models and datasets predominantly concentrate on accurately predicting the authenticity of news articles, often overlooking the determination of the reasons behind fake news and the interpretability of the results. To address this limitation, we propose a novel fine-grained classification task leveraging the fine-grained annotations provided in our dataset. This task extends the traditional binary classification of true or fake into a six-classification problem with fine-grained fake type annotation, including text-based fake, image-based fake, text-image consistency, content-knowledge inconsistency and others. In the training process, we utilize cross-entropy loss as loss function:
(2) $\mathcal{L}(y,\hat{y})=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i}),$
@@
Multi-domain adaptation task. Most existing fake news detection models are primarily trained within a specific domain, often overlooking the necessity to evaluate their effectiveness in accurately detecting fake news across multiple domains. However, considering the substantial presence of news spanning various domains on the internet, it becomes imperative for a model to possess effective multi-domain adaptation capabilities.
To assess the capacity of existing models for multi-domain adaptation, we further design three unsupervised domain adaptation tasks: topic adaptation, platform adaptation, and dual domain adaptation. Topic adaptation is proposed to measure the models’ capability to overcome data distribution variance $P(x_{s})\neq P(x_{t})$ between different topics from the same platform. In this task, models are trained using samples from four topics and then tested on the remaining two topics from the same platform. Platform adaptation, on the contrary, models are trained using samples from one platform while test on another platform with the same topic. This experiment aims to test the model’s ability to adapt to label shift problem $P(y_{s})\neq P(y_{t})$. Here, models are trained using samples from one platform and subsequently tested on another platform with the same topic. The most challenging task, dual domain adaptation, requires models to simultaneously adapting across both topics and platforms. This task presents a significant challenge, as it requires the model to exhibit generalization abilities enabling it to perform well when confronted with data from previously unseen domains and deviated label distribution.
4. Comprehensive Analysis of FineFake
Table 4. The numbers of sample distributions, average words and average entities in different domains of FineFake
@@ 4. Comprehensive Analysis of FineFake
In this section, we conduct an extensive experimental analysis aimed at unraveling the intricate structure and inherent characteristics of the FineFake dataset. Table 4 exhibits the detailed statistics of FineFake, including the number of true and fake samples in each topic/platform, the average text length, and the average number of extracted entities for each news. Upon scrutiny of these findings, it becomes apparent that our dataset exhibits a well-balanced distribution between positive and negative samples overall, ensuring a robust representation of both true and fake instances. Furthermore, the obvious differences in text length across three platforms (Stream Media, Official News, Fact-Check Websites) ensures the diversity of data. The imbalanced distribution of positive and negative samples in the three platforms are also significant. Official news sources generally exhibit higher credibility, thus having a higher proportion of true news. In contrast, when it comes to fact-checking websites, due to their focus on debunking fake news events, it has a higher proportion of fake news. This results not only aligns with our real-world understanding but also provides a solid data foundation for studying cross-platform label shift issues in fake news detection. Additionally, Table 4 demonstrates a substantial presence of annotated entity knowledge, affirming the richness and depth of contextual information within the dataset.
@@
Based on the empirical exploration, Figure 2(a) presents the top-30 words observed in the six topics and three platforms. Notably, one can easily observe that high-frequency vocabulary in different domains exhibits distinct patterns and thematic clusters. Such findings shed light on the domain-specific linguistic variations, revealing the importance of constructing a multi-domain fake news dataset. In order to delve deeper into the disparities in emotions between fake news and real news across domains, we conduct another experiment in terms of emotion tendency. Figure 2(b) and (c) are heatmaps illustrate the average value of emotion tendencies of the nine domains, which are calculated by Emotion DistilRoBERTa (hartmann2022emotionenglish, ). The emotion distributions of various domains are apparently different, further demonstrating the value of FineFake dataset. One can also see that the emotion tendencies of true/fake news are also different within the same domain, such as the “society” topic exhibits a more prominent “disgust” emotion in fake news than the real ones. In ”conflict” domain, there is a significant decrease in neutral sentiment compared to other domains, replaced by a predominant sense of anger. This finding is aligned with previous literature (alonso2021sentiment, ; bhutani2019fake, ) that sentiments also contribute to advancing the performance of fake news detection.
In order to scrutinize the presence of external knowledge entities, we conduct an analysis on the distribution of the number of external knowledge entities involved, as depicted in Figure 2(f). The results reveal that approximately 85% of the news articles encompass external knowledge entities, with real news articles demonstrating a higher propensity to incorporate such entities compared to fake news articles. Figure 2(d) illustrates the top-10 most frequently extracted entities and the top-3 entities are “Twitter”, “Donald Trump” and “USA President”. Figure 2(e) provides an overview of the proportions of data with fine-grained labels. This demonstrates FineFake’s exploration into the fundamental causes of fake news, thus providing a reliable benchmark for fine-grained analysis in future work. Moreover, the proportions of ""text-image inconsistency"" and ""content-knowledge inconsistency"" within the fake news category also highlight the importance of multimodal information and external knowledge. Figure 2(g) provides an overview of the proportions of data from each topic domain in relation to the total dataset, thereby highlighting the multi-domain characteristic of our study.
5. KEAN: The Proposed Method
5.1. Architecture
Figure 3. The overview of KEAN model.
@@ 5.1. Architecture
To address both the covariate shift and label shift problem simultaneously, we propose a knowledge-enhanced domain adaptation network, dubbed KEAN. As Fig 3 illustrates, the structure of KEAN is based on the architecture of Domain Adversarial Neural Networks (DANN) (ganin2016domain, ). As knowledge graph contains rich domain commonsense knowledge (chen2023knowledge, ; zeng2023knowledge, ; ghosal2020kingdom, ), we utilize the extracted entities to establish the sub-graph, thereby aiding in mitigating the covariate shift problem. Furthermore, inspired by previous work (garg2023rlsbench, ; li2019target, ; tachet2020domain, ), we incorporate a re-weighting strategy to solve label shift problem. Specifically, we re-weight the classifier by estimating the target label marginal to calculate label weights, thereby adjusting the influence of different labels on the classifier during training.
5.2. Multimodal Encoder
@@
We define an instance $I=\{T,V,E_{T}^{n}\}$ as a tuple representing three different modalities of contents: the textual content $T$ and the visual content $V$ of the news, and a set of textual entities $E_{T}^{n}$, where $n$ represents the number of entities. Additionally, we define the constructed knowledge graph is defined as $KG$.
Content Encoder. Considering both texutal content $T$ and visual content $V$ of the news, we utilize a pre-trained CLIP (Radford2021LearningTV, ) as the encoder in our methodology. This model is utilized to transform both sentences and images into embedding vectors, leveraging its robust multimodal representation capabilities (zhang2022hierarchical, ).
@@
Knowledge Graph Encoder. In FineFake, each instance contains accurate and abundant external knowledge information from Wikidata. We extract one-hop neighbours of the entities $E_{T}^{n}$ and construct a sub-graph of $KG$ by aggregating all the triplets, which is defined as $KG_{sub}$. Specifically, we utilize TransE (bordes2013translating, ) as our knowledge graph embedding method due to its simplicity and effectiveness. Formally, given a triplet $(h,r,t)$, TransE treats the relationship $r$ as a translation vector $r$ from the head entity $h$ to the tail entity $t$, represented as $h+r=t$. After feature extraction, each node $j\in KG_{sub}$ has its representation $h_{j}$. We then compute the average of the feature vectors $h_{j}$ for all nodes in $KG_{sub}$ to obtain the final graph feature $h_{kg}$ as the representation.
(4) $h_{j}=\mathrm{TransE}(E_{j}),E_{j}\in KG_{sub}$
$h_{kg}=\frac{1}{|E|}\sum_{E_{j}}^{E}h_{j},h_{j}\in KG_{sub}$
$h_{kg}^{{}^{\prime}}=\mathrm{MLP}_{kg}(h_{kg}),h_{kg}^{{}^{\prime}}\in\mathcal{R}^{d3}$
As data has encoded by pretrained model, we utilize fully connected layers for $h_{t},h_{v},h_{kg}$ to get the final representation $h_{t}^{{}^{\prime}},h_{v}^{{}^{\prime}},h_{kg}^{{}^{\prime}}$. In the fusion module, we use concatenation to get the final feature representation $h_{I}=[h_{t}^{{}^{\prime}};h_{v}^{{}^{\prime}};h_{kg}^{{}^{\prime}}]$ of the news I.
Table 5. Classification task results, best results are in bold and second best results are underlined.
@@
5.3. Adversarial Training Scheme
Domain-adversarial Training. To enable the model to learn domain-invariant representation and address covariate shift, we follow the architecture of DANN. KEAN comprises a task classifer $C$ (with parameters $\theta_{C}$), a domain-discriminator $D_{adv}$ (with parameters $\theta_{D}$) and a decoder $D_{recon}$ (with parameters $\theta_{R}$). $C$ and $D_{adv}$ both take fused representation $h_{I}$ as its inputs. The task classifier $C$ is responsible for determining the veracity of the input news $I$, while the domain-discriminator $D_{adv}$ aims to discriminate the domain label of $I$. The loss function are as followed:
(5) $\mathcal{L}_{cls}=E_{I_{s}}$ $(-\sum_{i=1}^{N}y_{i}\log C(h_{I}))$
To further enforce domain-invariance into the encoded representation $h_{kg}^{{}^{\prime}}$, we utilize a decoder $D_{recon}$ with a reconstruction loss, as proposed in prior work (bousmalis2016domain, ):
(6) $\mathcal{L}_{recon}(I_{s},I_{t})=E_{h_{kg}}(\left\|D_{recon}(h_{kg}^{{}^{\prime}})-h_{kg}\right\|^{2})$
The final optimization of the domain-adversarial training is based on the minimax game: where $\alpha$ and $\beta$ are hyper-parameters. The minimax game is realized by reversing the gradients of $\mathcal{L}_{adv}$ while back-propagation with a reverse layer (ganin2016domain, ).
(7) $\mathcal{L}_{loss}=(\mathcal{L}_{cls}+$ $\alpha\mathcal{L}_{adv}+\beta\mathcal{L}_{recon})$
Re-weighting. Re-weighting the classifier is a widely used technique to address label shift problems (garg2023rlsbench, ). Following the approach outlined in BBSE (lipton2018detecting, ), we aim to reweight the classifier by estimating the distribution of labels in target domain. Specifically, We seek to obtain the importance weights $\hat{w}_{t}(y)$, defined as the ratio of the probability of observing label $y$ in the target domain to that in the source domain, i.e., $\hat{w}_{t}(y)=\frac{p_{t}(y)}{p_{s}(y)}$. To calculate $\hat{w}_{t}(y)$, we first train the classifier and then calculate the confusion matrix $C_{h}$ of the source domain and probability mass function $q_{h}$ of $f(X)$ under predicted target distribution. Then, $\hat{w}_{t}(y)$ can be calculated by $C_{h}$ and $q_{h}$ as Equition 8. As we obtain $\hat{w}_{t}(y)$, we retrain the classifier by the importance weighted loss as followed:
(8) $\hat{w}_{t}(y)=$ $C_{h}^{-1}*q_{h}$
6. Experiments
6.1. Baseline Models

### v3

@@ III-A Multi-domain News Collection
Given our overarching objective of constructing a multimodal dataset and benchmark exclusively in English, text-only posts and news articles published in languages other than English are filtered. Subsequently, data deduplication techniques is implemented to eradicate redundant news items, thereby mitigating the risk of data leakage. Moreover, to uphold data quality standards, we eliminate instances containing unqualified images or excessively brief textual content.
Ultimately, instances are collected from three different types of platforms among six topics by using Snopes as the base, ensuring the diversity of FineFake. Each data point includes textual content, corresponding images, metadata, and any associated social media presence. The integration of multiple platforms and diverse topics underscores the multi-domain characteristics of FineFake, laying the groundwork for subsequent cross-platform and cross-topic research.
III-B Semi-manual Knowledge Alignment
@@
One notable feature of FineFake is its incorporation of pertinent background knowledge, which complements the original multi-modal content and holds immense potential to advance the identification of fake news. A semi-manual knowledge alignment strategy is adopted to accomplish the integration. Specifically, an entity link tool  [39] is employed to adeptly recognize named entities within the news text. To ensure the accuracy of extracted named entities, we initially set a lower threshold $\omega_{1}$ to compile the initial list of entities and subsequently adjust the threshold higher to $\omega_{2}$ to identify entities with higher confidence. Entities extracted within the threshold range of $\omega_{1}$ to $\omega_{2}$ are verified by human annotators for contextual accuracy, achieving disambiguation. Subsequently, all triplets within Wikidata with a distance of one from the entity are retrieved, including relationship-type triplets and attribute triplets. Relationship-type triplets denotes as $(h,r,t)$, where $h$ denotes the head entity, $r$ denotes the relation and $t$ denotes the tail entity. They represent entities connected to the extracted entities in Wikipedia, these triplets are widely utilized by knowledge-enhanced methods to capture the background knowledge of news [40, 16]. Attribute triplets denotes as $(e,r_{d},d)$, where $e$ denotes the entity, $r_{d}$ denotes the relation is description of the entity and $d$ denotes the text of description. This augmentation adds knowledge assistance, thereby facilitating more nuanced analyses and interpretations of fake news instances.
III-C Fine-grained Human Annotation
@@
Therefore, in the verification stage, 5 professional annotators are engaged to annotate the binary classification label and fine-grained label for each data with the usage of Snopes as references. To promote consistency and mitigate subjectivity, we implement inter-annotator agreement measures. Each news is independently labeled by five annotators to assess consistency and calculate agreement scores (e.g. Fleiss’s kappa). Discrepancies are resolved through discussions among annotators. The refined categorization scheme in FineFake facilitates a deeper analysis of the underlying reasons of fake news, empowering researchers to develop sophisticated detection methods that account for distinct modalities, inconsistencies, and the intersection between textual and visual elements.
III-D Fine-Grained and Multi-Domain Fake News Detection Tasks
Based on the constructed FineFake dataset, we propose three downstream tasks to evaluate the performance of SOTA fake news detection models under various scenarios.
III-D1 Binary classification task
The primary objective in fake news detection models is to classify news articles as either fake or true. To investigate the efficacy of external knowledge in enhancing fake news detection, we conduct two types of binary classification tasks: one without any external knowledge and another with knowledge augmentation.
$\mathcal{L}(y,\hat{y})=-y\log(\hat{y})-(1-y)\log(1-\hat{y}),$ (1)
@@ III-D1 Binary classification task
Figure 3: Basic information and statistic analysis on FineFake.
III-D2 Fine-grained classification task
Currently, both fake news detection models and datasets predominantly concentrate on accurately predicting the authenticity of news articles, often overlooking the determination of the reasons behind fake news. To address this limitation, a fine-grained classification task is proposed to expand traditional binary classes into six classes, as detailed defined in Section III-C. In the training process, we utilize cross-entropy loss as loss function:
$\mathcal{L}(y,\hat{y})=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i}),$ (2)
@@ III-D2 Fine-grained classification task
III-D3 Multi-domain adaptation task
To evaluate cross-domain capacity of existing models, we design three cross-domain tasks: topic adaptation, platform adaptation and dual domain adaptation. Topic adaptation is proposed to measure the model’s capability to overcome data distribution variance between different topics from the same platform. In this task, models are trained on four topics and tested on the remaining two topics. Platform adaptation, on the contrary, aims to test the model’s ability to adapt to label shift problem, with training on one platform and testing on another under the same topic. The most challenging task, dual domain adaptation, requires models to simultaneously adapting across both topics and platforms. This task presents a significant challenge, as it requires the model to exhibit generalization abilities enabling it to perform well when confronted with data from previously unseen domains and deviated label distribution.
TABLE IV: The numbers of sample distributions, average words and average entities in different domains of FineFake.
@@ IV Comprehensive Data Analysis of FineFake
In this section, we analyze the FineFake dataset to understand its complex structure and characteristics. Table IV shows detailed statistics, including the number of true and fake samples in each topic/platform, the average text length, and average entities for each news. The obvious differences in text length across three categories of platforms ensures the diversity of data. Furthermore, the imbalanced distribution of positive and negative samples in the three platforms are also significant. Official news sources generally exhibit higher credibility, thus having a higher proportion of true news. In contrast, Snopes has a higher proportion of fake news as it focuses on debunking fake news events.
@@
(a) Multi-Domain Analysis. Figure 3(a) presents the top-30 words observed in the six topics and three platforms. Notably, one can easily observe that high-frequency vocabulary in different domains exhibits distinct patterns and thematic clusters. Such findings shed light on the domain-specific linguistic variations, revealing the importance of constructing a multi-domain fake news dataset. (b) Emotion Tendency. Figure 3(b) and (c) illustrate the average value of emotion tendencies of the nine domains, which are calculated by Emotion DistilRoBERTa [41]. The emotion distributions of various domains are apparently different, further demonstrating the value of FineFake dataset. One can also see that the emotion tendencies of true/fake news are also different within the same domain, such as the “society” topic exhibits a more prominent “disgust” emotion in fake news than the real ones. In ”conflict” topic, there is a significant decrease in neutral sentiment compared to other topics, replaced by a predominant sense of anger. This finding is aligned with previous literature [42, 43] that sentiments also contribute to advancing the performance of fake news detection. (c) External Knowledge Analysis. Figure 3(d) and (f) analyze the presence of external knowledge entities. The results reveal that approximately 85% of the news articles encompass external knowledge entities, with real news articles demonstrating a higher propensity to incorporate such entities compared to fake news articles. Figure 3(d) illustrates the top-9 most frequently extracted entities and the top-3 entities are “Twitter”, “Donald Trump” and “USA President”. (d) Data Proportion Analysis. Figure 3(e) provides an overview of the proportions of data with fine-grained labels. This demonstrates FineFake’s exploration into the fundamental causes of fake news, thus providing a reliable benchmark for fine-grained analysis in future work. Moreover, the proportions of “text-image inconsistency” and “content-knowledge inconsistency” within the fake news category also highlight the importance of multimodal information and external knowledge. Figure 3(g) provides an overview of the proportions of data from each topic domain in relation to the total dataset, thereby highlighting the multi-domain characteristic of our study.
Figure 4: The overview of KEAN model.
V KEAN: The Knowledge-Enhanced Domain Adaptation Network
@@
To address both the covariate shift and label shift problem simultaneously, we propose a knowledge-enhanced domain adaptation network, dubbed KEAN. As Fig 4 illustrates, the structure of KEAN is based on the architecture of Domain Adversarial Neural Networks (DANN) [44]. As knowledge graph contains rich domain commonsense knowledge [45, 46, 47], three encoders are utilized for text, visual and knowledge modeling, respectively. Furthermore, inspired by previous work [48, 49, 50], a reconstruction module and a reweight module is implemented to help solve the covariate shift and label shift simultaneously.
@@
Multimodal Encoder. An instance is defined as a tuple $I=\{T,V,E_{T}^{n}\}$ representing three modalities of contents: the textual content $T$ and the visual content $V$ of the news, and a set of textual entities $E_{T}^{n}$, where $n$ represents the number of entities. Additionally, the constructed knowledge graph is defined as $KG$. A pre-trained CLIP [51] is utilized as the encoder in our method, which transforms sentences and images into embeddings through its robust multimodal representation capabilities:
@@
Knowledge Graph Encoder. As each instance in FineFake contains external knowledge from wikidata, the one-hop neighbours of the entities $E_{T}^{n}$ are extracted to construct a sub-graph of $KG$ by aggregating all the triplets, defined as $KG_{sub}$. Specifically, TransE [52] is utilized as our knowledge graph embedding method due to its simplicity and effectiveness. After feature extraction, each node $j\in KG_{sub}$ has its representation $h_{j}$. Then the average of the feature vectors $h_{j}$ for all nodes in $KG_{sub}$ is computed to obtain the final graph feature $h_{kg}$ as the representation:
$h_{kg}=\frac{1}{|E|}\sum_{E_{j}}^{E}\mathrm{TransE}(E_{j}),E_{j}\in KG_{sub}$ (4)
The representations $h_{t},h_{v},h_{kg}$ are then transformed to $h_{t}^{{}^{\prime}},h_{v}^{{}^{\prime}},h_{kg}^{{}^{\prime}}$ through fully connected layers. The final feature representation is obtained by concatenation: $h_{I}=[h_{t}^{{}^{\prime}};h_{v}^{{}^{\prime}};h_{kg}^{{}^{\prime}}]$.
Domain-adversarial Training. Based on the DANN architecture, KEAN comprises a task classifer $C$ (with parameters $\theta_{C}$), a domain-discriminator $D_{adv}$ (with parameters $\theta_{D}$) and a decoder $D_{recon}$ (with parameters $\theta_{R}$). $C$ and $D_{adv}$ focus respectively on news truthfulness and domain differentiation:
$\mathcal{L}_{cls}=E_{I_{s}}$ $(-\sum_{i=1}^{N}y_{i}\log C(h_{I}))$ (5)
To further enforce domain-invariance into the encoded representation $h_{kg}^{{}^{\prime}}$, A decoder $D_{recon}$ is utilized with a reconstruction loss, as proposed in prior work [53]:
$\mathcal{L}_{recon}(I_{s},I_{t})=E_{h_{kg}}(\left\|D_{recon}(h_{kg}^{{}^{\prime}})-h_{kg}\right\|^{2})$ (6)
The final optimization of the domain-adversarial training is based on the minimax game: where $\alpha$ and $\beta$ are hyper-parameters. The minimax game is realized by reversing the gradients of $\mathcal{L}_{adv}$ while back-propagation with a reverse layer [44]:
$\mathcal{L}_{loss}=(\mathcal{L}_{cls}+$ $\alpha\mathcal{L}_{adv}+\beta\mathcal{L}_{recon})$ (7)
Re-weighting. Re-weighting the classifier is a widely used technique to address label shift problems [48]. Following the approach outlined in BBSE [54], the classifier is re-weighted by estimating the distribution of labels in target domain. Specifically, We seek to obtain the importance weights $\hat{w}_{t}(y)$, defined as the ratio of the probability of observing label $y$ in the target domain to that in the source domain, i.e., $\hat{w}_{t}(y)=\frac{p_{t}(y)}{p_{s}(y)}$. To calculate $\hat{w}_{t}(y)$, the classifier is trained with source data while the confusion matrix $C_{h}$ of the source domain and probability mass function $q_{h}$ of $f(X)$ under predicted target distribution is then calculated. Then, $\hat{w}_{t}(y)$ can be obtained by $C_{h}$ and $q_{h}$ as Equation 8. As $\hat{w}_{t}(y)$ is obtained, the classifier is retrained by the importance weighted loss as followed:
$\hat{w}_{t}(y)=$ $C_{h}^{-1}*q_{h}$ (8)
TABLE V: Classification task results, best results are in bold and second best results are underlined.
@@
VI Experiments
VI-A Baseline Models